OpenStudy (anonymous):

HELP ME ? ILL MEDAL AND FAN Which of the following is equivalent to the expression below? (9 + 2x) - 2x A. 9 B. 9 + 4x C. 9 - 4x D. 2

OpenStudy (mathstudent55):

In the given expression, the parentheses are not needed, so just drop them. Then, combine like terms.

OpenStudy (anonymous):

no A

OpenStudy (anonymous):

its A

OpenStudy (mathstudent55):

No. First, just drop the parentheses. You get \(9+ 2x - 2x\) Now look at 2x - 2x. They are like terms. If you subtract something (2x) from itself (2x) you get zero. Just like 10 - 10 = 0, 5 - 5 = 0, x - x = 0, also 2x - 2x = 0. That means the 2x - 2x part is zero, and you are just left with 9. Here is how this porblem looks like step by step: \((9 + 2x) - 2x\) \(= 9 + 2x - 2x \) \(= 9 + 0\) \(=9\)